OpenStudy (anonymous):

hi everyone I'll help just I do have rules RULES: 1. Please put your question in appropriate words 2. Write your username under your question like this example: @iGreen 3. And I will attempt to answer if not I will get a friend to help please follow my rules thank you!

OpenStudy (anonymous):

@TynisaNTaylor please write your question

OpenStudy (anonymous):

You are offered two different sales jobs. The first company offers a straight commission of 4% of the sales. The second company offers a salary of $ 390 per week plus 1% of the sales. How much would you have to sell in a week in order for the straight commission offer to be at least as good? @TynisaNTaylor

OpenStudy (anonymous):

@UnkleRhaukus (◊��◊��◊��◊��) help her please

OpenStudy (unklerhaukus):

$_1 = S * 4% $_2 = 390 + S*1% for $_1 to be as good as $_2 $_1 = $_2 S * 4% = 390 + S * 1% = solve for S

OpenStudy (unklerhaukus):

Where $_1 is the income from the first offer $_2 is '' '' '' second offer and S is sales